# Simulating the Visual World with Artificial Intelligence: A Roadmap

### v1

@@ Abstract
The landscape of video generation is shifting, from a focus on generating visually appealing clips to building virtual environments that support interaction and maintain physical plausibility. These developments point toward the emergence of video foundation models that function not only as visual generators but also as implicit world models, models that simulate the physical dynamics, agent-environment interactions, and task planning that govern real or imagined worlds. This survey provides a systematic overview of this evolution, conceptualizing modern video foundation models as the combination of two core components: an implicit world model and a video renderer. The world model encodes structured knowledge about the world, including physical laws, interaction dynamics, and agent behavior. It serves as a latent simulation engine that enables coherent visual reasoning, long-term temporal consistency, and goal-driven planning. The video renderer transforms this latent simulation into realistic visual observations, effectively producing videos as a “window” into the simulated world. We trace the progression of video generation through four generations, in which the core capabilities advance step by step, ultimately culminating in a world model, built upon a video generation model, that embodies intrinsic physical plausibility, real-time multimodal interaction, and planning capabilities spanning multiple spatiotemporal scales. For each generation, we define its core characteristics, highlight representative works, and examine their application domains such as robotics, autonomous driving, and interactive gaming. Finally, we discuss open challenges and design principles for next-generation world models, including the role of agent intelligence in shaping and evaluating these systems. An up-to-date list of related works is maintained at this link.
@@ 1.1 Motivation
World models, which aim to simulate the real world, have long posed significant challenges in artificial intelligence, influencing various applications, such as robotics, autonomous driving, and gaming. While the specific capabilities required for world models are numerous and not yet precisely defined, approaches such as 3D generation [lumagenie2025, liu2025unleashing, huang2024placiddreamer], 3D/4D scene generation, and video generation demonstrate one or more relevant abilities, such as motion dynamics, interaction and controllability, visual quality, 3D consistency, and generation efficiency. Consequently, these approaches have been adopted in recent works as potential pathways toward realizing world models.
@@
Recent breakthroughs [assran2025vjepa2, mirage2025, klingai2024kling, sora, pika23, jimengai24jimeng, yang2024cogvideox, lin2025apt2, agarwal2025cosmos, alhaija2025cosmostransfer1, runway2025gen4, Veo325, chen2025skyreels, stepvideo25jieyue, gao2025seedance, bao2024vidu, luma24, yin2025causvid, hailuo0225minmax, teng2025magi, wan2025wan, kong2024hunyuanvideo, fan2025vchitect] in video generation models, driven by improvements in diffusion models [hatamizadeh2024diffit, zhang2023improved, song2023lossguided, mardani2023variational, song2023pseudoinverse, lim2023score, tashiro2021csdi, song2020ddim, zhu2025asdm, zhong2025vfrtok], autoregressive backbones [meng2020autoregressive], variational autoencoders [zhao2019infovae, zhao2017learning], image generation techniques [dang2025personalized, zhang2023diffcollage, ye2023affordance, balaji2022ediff, meng2021improvedar, meng2021sdedit, luma24, zhao2017towards, liu2025rpe, shen2024sgadapter], controllable image generation  [atzmon2024edifyimagediff, sinha2021d2c], and enhancements in training or inference efficiency [zhou2025inductive, song2025ideas, xu2024agg, ozturkler2023smrd, he2025scaling, wu2025vmoba, zhang2025Jenga, shi2025diffmoe, yang2025efficient, yin2025towards, han2024agent], as well as more flexible condition injection modules [zhang2025motioncrafter, zhang2023controlnet], and advances in video rendering [liu2025improving], mark a pivotal moment for the field. These models can now generate high-quality videos grounded in fundamental world knowledge. At this stage, they begin to play a central role in constructing world models. Moreover, with the rapid advancement of technologies such as virtual reality (VR) and embodied AI, the integration of world models into interactive, real-time environments has become increasingly feasible. These technological trends, together with the maturation of video generation methods, indicate that we are on the cusp of a new era where world models will play a central role in shaping autonomous systems, intelligent agents, and immersive virtual environments. This trend is further evidenced in Figure 2, which illustrates the dynamics of research attention over recent years. Specifically, while discussions and mentions of world models have been present since 2018, they have remained relatively steady in terms of annual publication volume. While starting in 2024, video generation witnessed an explosive surge in both technical advances and the number of related works, which in turn catalyzed a renewed wave of progress in world models. This pattern highlights a clear interdependence: the rapid advances in video generation are not merely parallel developments but are becoming key enablers that support and accelerate the evolution of world models. Consequently, the present moment is particularly critical for the community to systematically discuss the evolution from video generation to world models, and to chart their future directions as an integrated research frontier.
@@
Despite these advances, challenges remain at both the conceptual and structural levels. Conceptually, the definition of a “world model” remains ambiguous, making it difficult to unify perspectives and evaluate progress. Structurally, the field lacks a well-established taxonomy to organize modeling capabilities, developmental stages, and potential trajectories. These gaps highlight an urgent need for a systematic elucidation and comprehensive survey to consolidate existing knowledge and guide the next stage of research. Current surveys [wang2025survey, xing2024survey, yu2025surveyigv, dal2024jointadsurvey, zhu2024issoraasurvey, fu2024exploringsurvey, melnik2024videosurvey, lin2025exploringsurvey, guan2024adinitialsurvey, liu2025generativesurvey, cho2024soraasansgisurvey, sun2024fromsorawecanseesurvey, yu2025position] have laid important groundwork by summarizing related methodologies, datasets, and applications. However, there remains a need to explicitly clarify what aspects have been thoroughly addressed, and which areas, such as real-time integration, controllable video-to-world pipelines, or comprehensive evaluation metrics, are still underexplored. Accordingly, this survey aims to chart a clear path from video generation toward comprehensive world modeling, providing guidance for future research and development in this emerging field.
@@ 1.2 Position
From this standpoint, the physical world model represents a more fundamental and global conceptualization, aiming to capture both the evolution and the intrinsic laws of the physical world. In contrast, the mental world model [kessler2023effectiveness, xu2022learning, lu2022challenges, ball2021augmented] can be regarded as a specialized internal cognitive framework that may emerge in higher‑generation physical world models, serving to represent an agent’s internal states, intentions, and preferences.
@@ 3 Roadmap
Superficial Faithfulness. Faithfulness is the primary capability emphasized in Generation 1, as it represents the foundational step from video generation towards world models, which shifts from static realism to realistic motion dynamics. This generation is categorized into two stages: the Basic Level and the Advanced Level. The Basic Level consists of traditional and classical approaches, primarily based on GANs and early diffusion models. In contrast, the Advanced Level encompasses video generation models that exhibit superficial faithfulness and low-level interactiveness, while still lacking explicit planning capabilities.
In this axis, short and long faithfulness denote the temporal span over which realistic and physically consistent generations can be maintained. We adopt a practical threshold where long corresponds to sequences of five seconds or longer at 24 FPS, following the convention that such durations begin to reveal a model’s stability as a world simulator rather than a mere simple video generator or single motion synthesizer. This temporal coherence implicitly depends on the model’s internal memory mechanisms, even though memory itself is not presented as a separate capability axis. Thus, although memory is important capability to the operation of any world model, we treat it as an underlying mechanism that enhances along with the mentioned observable capabilities, such as temporal faithfulness and consistent planning rather than as a standalone dimension in this taxonomy.
@@
Consistent Faithfulness. Beyond improved interactiveness, Generation 2 world models show significant progress in maintaining temporal coherence and video–text consistency across long and complex sequences. The ability to generate longer videos contributes directly to temporal consistency, ensuring stable object dynamics and scene layouts over time. Moreover, these models achieve perfect video-text consistency, faithfully rendering all mentioned entities, motions, and events aligned with the given inputs. This generation also begins to capture aspects of the basic physical world, including projective geometry and spatial appropriateness, which contribute to more physically plausible and coherent video generations.
@@ 4 Methods: A Hierarchical Taxonomy
Open-Sora 2.0 [peng2025opensora2] DiT 120 $1024\times 576$ SkyReels-V2 [chen2025skyreels] DiT 121 $1280\times 720$
@@
Genie 3 [google2025genie3] AR >1440>1440 $1280\times 720$
@@ 4.1.1 Foundation Models with Visual-Centric World Knowledge
We summarize representative pre-trained text-to-video (T2V) models [hong2022cogvideo, chen2023videocrafter1, chen2024videocrafter2, blattmann2023svd, wang2023modelscope, gen223, guo2023animatediff, zhang2024show1, wang2025lavie, qing2024higen, ma2024latte, wang2024magicvideov2, bar2024lumiere, klingai2024kling, henschel2025streamingt2v, girdhar2023emuvideo, kondratyuk2023videopoet, yuan2024instructvideo, zhou2024allegro], along with commonly used pre-trained image-to-video (I2V) models [zhang2023i2vgen, lin2024opensoraplan, hailuo24minmax, xu2024easyanimate, klingai2024kling, blattmann2023svd, yu2023magvitv2, chen2023videocrafter1, sun2024dimensionx, kondratyuk2023videopoet, guo2024i2vadapter], examining their backbone architectures, video lengths, resolutions, and post-training capabilities related to interactivity and embodiment which are core abilities for serving as world models. This survey also covers both open-sourced models [ma2024latte, wang2025lavie, xu2024easyanimate, si2025repvideo, chen2024videocrafter2, guo2023animatediff, hacohen2024ltxvideo, chen2023videocrafter1, zheng2024opensora, zhang2024show1, lin2024opensoraplan, qing2024higen, wang2023modelscope, henschel2025streamingt2v, sun2024dimensionx, kondratyuk2023videopoet, yuan2024instructvideo, zhou2024allegro] and commercial APIs [hailuo24minmax, klingai2024kling, gen223, pika23, girdhar2023emuvideo]. For instance, CogVideo [hong2022cogvideo] is widely regarded as an early open-source, large-scale pretrained T2V model that played a pivotal role in initiating the development of video foundation models. The release of Kling 1.0 [klingai2024kling], built on a DiT backbone, further advanced generative AI toward longer duration and greater visual quality.
@@ 4.2.1 Foundation Models with Dynamic-Aware World Knowledge
Recent research builds on this momentum, consolidating video generation as the tool for controllable world models. These models [wan2025wan, peng2025opensora2, genmo2024mochi1, xu2024easyanimate, gao2025seedance, chen2025skyreels, kong2024hunyuanvideo, sora, yang2024cogvideox, klingai2024kling, imaginegrok, wan22, openai2025sora2, luma2025ray3] commonly adopt DiT-based [peebles2023dit] or some of them leverage [lin2025apt2, yin2025causvid, huang2025selfforcing] hybrid architectures that combine diffusion models with causal, frame-by-frame autoregressive mechanisms, achieving a fast, few-step distilled diffusion process to generate each frame. For instance, models such as MAGI‑1 [teng2025magi] and CausVid [yin2025causvid] extend Diffusion Forcing [chen2024diffusionforcing] for causal video generation. Self Forcing [huang2025selfforcing] follows this paradigm but removes exposure bias by training with full autoregressive rollout that matches the true inference distribution, enabling efficient video‑level supervision and semi-real‑time generation. And VMoBA [wu2025vmoba] introduces a mixture-of-block attention mechanism to enhance spatiotemporal representation learning in video diffusion models, achieving improved generation quality and efficiency.
@@ 4.2.2 Video Generation as World Model in General Scene
In this section, we introduce methods that transition from video generation to world modeling in general-purpose scenarios, with representative methods summarized in Table 3. The term general follows its definition in the broader video generation, referring to settings that are not tailored to any specific downstream application. Our discussion focuses on two main groups of methods: (1) Approaches with spatial condition [sketchvideo, das, bian2025gsdit, li2025pisa, xue2025phyt2v, wang2025wisa], which incorporate various priors such as Geometry Conditions, 3D Priors, and Physical Priors. (2) Approaches with navigation mode [fu20243dtrajmaster, he2025cameractrl2, ren2025gen3c, hou2024camtrol, aether, hong2024slowfast, xiang2024pandora], which condition generation on forms of controllable input, including action, trajectory, camera motion, and text instruction.
@@
Physical Prior World Model. Physical prior world models [li2025pisa, wang2025wisa, xue2025phyt2v, zhao2025synthetic] aim to enhance the physical plausibility of video generation by embedding explicit physical laws and principles into the modeling process. For instance, WISA [wang2025wisa] incorporates physical formulas and principles directly as embeddings. PhyT2V [xue2025phyt2v] leverages the physical reasoning capabilities of large language models (LLMs) to guide and improve the physical consistency of video generation models.
@@
Instruction Navigation World Model. Instruction Navigation World Models [aether, xiang2024pandora, hong2024slowfast] utilize text-based instructions, but unlike conventional text prompts that describe background elements, such as scene type, weather, or task themes, these instructions instead guide the dynamic aspects of video generation. Specifically, they control motion-related factors of either a single or multiple subjects in the video, or the perspective dynamics of a first-person agent, which can be a human, robot, or camera. For instance, models like Pandora [xiang2024pandora] and SlowFast-Gen [hong2024slowfast] can execute directional commands such as “turn left” or “turn right.” In addition, SlowFast-Gen [hong2024slowfast] supports perspective-based actions such as zooming in and out from a navigational viewpoint.

### v2

@@ Abstract
The landscape of video generation is shifting, from a focus on generating visually appealing clips to building virtual environments that support interaction and maintain physical plausibility. These developments point toward the emergence of video foundation models that function not only as visual generators but also as implicit world models, models that simulate the physical dynamics, agent-environment interactions, and task planning that govern real or imagined worlds. This survey provides a systematic overview of this evolution, conceptualizing modern video foundation models as the combination of two core components: an implicit world model and a video renderer. The world model encodes structured knowledge about the world, including physical laws, interaction dynamics, and agent behavior. It serves as a latent simulation engine that enables coherent visual reasoning, long-term temporal consistency, and goal-driven planning. The video renderer transforms this latent simulation into realistic visual observations, effectively producing videos as a “window” into the simulated world. We trace the progression of video generation through four generations, in which the core capabilities advance step by step, ultimately culminating in a world model, built upon a video generation model, that embodies intrinsic physical plausibility, real-time multimodal interaction, and planning capabilities spanning multiple spatiotemporal scales. For each generation, we define its core characteristics, highlight representative works, and examine their application domains such as robotics, autonomous driving, and interactive gaming. Finally, we discuss open challenges and design principles for next-generation world models, including the role of agent intelligence in shaping and evaluating these systems.
@@ 1.1 Motivation
World models, which aim to simulate the real world, have long posed significant challenges in artificial intelligence, influencing various applications, such as robotics, autonomous driving, and gaming. While the specific capabilities required for world models are numerous and not yet precisely defined, approaches such as 3D generation [lumagenie2025, liu2025unleashing, huang2024placiddreamer, yang2019pointflow, cao2025physx], 3D/4D scene generation [sun20253d, hsu2025programs, yang2024holodeck, yu2025wonderworld, yu2024wonderjourney, huang2025omnix, huang2025terra, zhaoxi20254dnex], and video generation demonstrate one or more relevant abilities, such as motion dynamics, interaction and controllability, visual quality, 3D consistency, and generation efficiency. Consequently, these approaches have been adopted in recent works as potential pathways toward realizing world models.
@@
Recent breakthroughs [assran2025vjepa2, mirage2025, klingai2024kling, sora, pika23, jimengai24jimeng, yang2024cogvideox, lin2025apt2, agarwal2025cosmos, alhaija2025cosmostransfer1, runway2025gen4, Veo325, chen2025skyreels, stepvideo25jieyue, gao2025seedance, bao2024vidu, luma24, yin2025causvid, hailuo0225minmax, teng2025magi, wan2025wan, kong2024hunyuanvideo, fan2025vchitect] in video generation models, driven by improvements in diffusion models [hatamizadeh2024diffit, zhang2023improved, song2023lossguided, mardani2023variational, song2023pseudoinverse, lim2023score, tashiro2021csdi, song2020ddim, zhu2025asdm, zhong2025vfrtok, shi2025latent, zhong2025decoupling], autoregressive backbones [meng2020autoregressive], variational autoencoders [zhao2019infovae, zhao2017learning], image generation techniques [rombach2022high, dang2025personalized, zhang2023diffcollage, ye2023affordance, balaji2022ediff, meng2021improvedar, meng2021sdedit, luma24, zhao2017towards, liu2025rpe, shen2024sgadapter, zeng2024jedi, ho2020denoising, si2024freeu, huang2024reversion, shi2025svg, xian2025free, kumbong2025hmar, gordon2025stencil], controllable image generation  [zhang2023controlnet, atzmon2024edifyimagediff, sinha2021d2c, mou2024t2i, huang2023collaborative, wang2023styleadapter], and enhancements in training or inference efficiency [zhou2025inductive, song2025ideas, xu2024agg, ozturkler2023smrd, he2025scaling, wu2025vmoba, zhang2025Jenga, shi2025diffmoe, yang2025efficient, yin2025towards, han2024agent, ju2025fulldit, he2025fulldit2, zhang2025training, zhengyao2025dcm], as well as more flexible condition injection modules [zhang2025motioncrafter, zhang2023controlnet], mark a pivotal moment for the field. To this end, recent progress further expands the capabilities of video generation across video reasoning and alignment [huang2025vchain, liu2025improving, li2025gran, wang2025monet, yin2025sea, cheng2025realdpo, yuan2024instructvideo], long-horizon storytelling and film generation [wang2025mavis, he2024storytelling, huang2025filmaster, zhuang2024vlogger, zhao2024moviedreamer, xie2024dreamfactory, wu2025automated, he2025cut2next], high-resolution video generation [qiu2025cinescale, he2024venhancer], personalized or multi-concept video customization [huang2025conceptmaster, wu2025customcrafter], and test-time scaling (TTS) [he2025scaling]. Collectively, these advances reinforce the emerging view that modern video generation systems are becoming fundamental components for constructing world models. Moreover, with the rapid advancement of technologies such as virtual reality (VR) and embodied AI, the integration of world models into interactive, real-time environments has become increasingly feasible. These technological trends, together with the maturation of video generation methods, indicate that we are on the cusp of a new era where world models will play a central role in shaping autonomous systems, intelligent agents, and immersive virtual environments. This trend is further evidenced in Figure 2, which illustrates the dynamics of research attention over recent years. Specifically, while discussions and mentions of world models have been present since 2018, they have remained relatively steady in terms of annual publication volume. While starting in 2024, video generation witnessed an explosive surge in both technical advances and the number of related works, which in turn catalyzed a renewed wave of progress in world models. This pattern highlights a clear interdependence: the rapid advances in video generation are not merely parallel developments but are becoming key enablers that support and accelerate the evolution of world models. Consequently, the present moment is particularly critical for the community to systematically discuss the evolution from video generation to world models, and to chart their future directions as an integrated research frontier.
@@
Despite these advances, challenges remain at both the conceptual and structural levels. Conceptually, the definition of a “world model” remains ambiguous, making it difficult to unify perspectives and evaluate progress. Structurally, the field lacks a well-established taxonomy to organize modeling capabilities, developmental stages, and potential trajectories. These gaps highlight an urgent need for a systematic elucidation and comprehensive survey to consolidate existing knowledge and guide the next stage of research. Current surveys and benchmarks [kong20253d, wang2025survey, xing2024survey, yu2025surveyigv, dal2024jointadsurvey, zhu2024issoraasurvey, fu2024exploringsurvey, melnik2024videosurvey, lin2025exploringsurvey, guan2024adinitialsurvey, liu2025generativesurvey, cho2024soraasansgisurvey, sun2024fromsorawecanseesurvey, yu2025position, duan2025worldscore, qin2024worldsimbench, li2025worldmodelbench, zheng2025vbench2, zhang2025evaluationagent] have laid important groundwork by summarizing related methodologies, datasets, and applications. However, there remains a need to explicitly clarify what aspects have been thoroughly addressed, and which areas, such as real-time integration, controllable video-to-world pipelines, or comprehensive evaluation metrics, are still underexplored. Accordingly, this survey aims to chart a clear path from video generation toward comprehensive world modeling, providing guidance for future research and development in this emerging field.
@@ 1.2 Position
From this standpoint, the physical world model represents a more fundamental and global conceptualization, aiming to capture both the evolution and the intrinsic laws of the physical world. In contrast, the mental world model [kessler2023effectiveness, xu2022learning, lu2022challenges, ball2021augmented, kolev2024efficient] can be regarded as a specialized internal cognitive framework that may emerge in higher‑generation physical world models, serving to represent an agent’s internal states, intentions, and preferences.
@@ 3 Roadmap
Superficial Faithfulness. Faithfulness is the primary capability emphasized in Generation 1, as it represents the foundational step from video generation towards world models, which shifts from static realism to realistic motion dynamics. This generation is categorized into two stages: the Basic Level and the Advanced Level. The Basic Level consists of traditional and classical approaches, primarily based on GANs [liu2021generative] and early diffusion models. In contrast, the Advanced Level encompasses video generation models that exhibit superficial faithfulness and low-level interactiveness, while still lacking explicit planning capabilities.
In this axis, short and long faithfulness denote the temporal span over which realistic and physically consistent generations can be maintained. We adopt a practical threshold where long corresponds to sequences of five seconds or longer at 24 FPS, following the convention that such durations begin to reveal a model’s stability as a world simulator rather than a mere simple video generator or single motion synthesizer. This temporal coherence implicitly depends on the model’s internal memory mechanisms, even though memory itself is not presented as a separate capability axis. Thus, although memory [po2025long] is important capability to the operation of any world model, we treat it as an underlying mechanism that enhances along with the mentioned observable capabilities, such as temporal faithfulness and consistent planning rather than as a standalone dimension in this taxonomy.
@@
Consistent Faithfulness. Beyond improved interactiveness, Generation 2 world models show significant progress in maintaining temporal coherence and video–text consistency across long and complex sequences. The ability to generate longer videos contributes directly to temporal consistency [wu2024freeinit], ensuring stable object dynamics and scene layouts over time. Moreover, these models achieve perfect video-text consistency, faithfully rendering all mentioned entities, motions, and events aligned with the given inputs. This generation also begins to capture aspects of the basic physical world, including projective geometry and spatial appropriateness, which contribute to more physically plausible and coherent video generations.
@@ 4 Methods: A Hierarchical Taxonomy
Open-Sora 2.0 [peng2025opensora2] DiT 120 $1024\times 576$ SkyReels-V2 [chen2025skyreels] AR-DiT 121 $1280\times 720$
@@
Kling-Omni [team2025kling] DiT $240$ $1920\times 1080$ SemanticGen [bai2025semanticgen] DiT $1440$ $832\times 480$
MemFlow [ji2025memflow] AR-DiT $1122$ $832\times 480$
@@
Genie 3 [google2025genie3] AR >1440>1440 $1280\times 720$ Magica 2 [magica22025] AR-DiT Minutes -
@@
PhysDreamer [zhang2024physdreamer] 3D Action Video Generation -
WonderPlay [li2025wonderplay] None Action Video Generation Go-with-the-Flow [burgert2025gowiththeflow]
@@
MotionStream [shin2025motionstream] None Trajectory Video Generation Wan Series [wan2025wan, wan22]
@@ 4.1.1 Foundation Models with Visual-Centric World Knowledge
We summarize representative pre-trained text-to-video (T2V) models [hong2022cogvideo, chen2023videocrafter1, chen2024videocrafter2, blattmann2023svd, wang2023modelscope, gen223, guo2023animatediff, zhang2024show1, wang2025lavie, qing2024higen, ma2024latte, wang2024magicvideov2, bar2024lumiere, klingai2024kling, henschel2025streamingt2v, girdhar2023emuvideo, kondratyuk2023videopoet, yuan2024instructvideo, zhou2024allegro, tian2024videotetris], along with commonly used pre-trained image-to-video (I2V) models [zhang2023i2vgen, lin2024opensoraplan, hailuo24minmax, xu2024easyanimate, klingai2024kling, blattmann2023svd, yu2023magvitv2, chen2023videocrafter1, sun2024dimensionx, kondratyuk2023videopoet, guo2024i2vadapter, team2025kling], examining their backbone architectures, video lengths, resolutions, and post-training capabilities related to interactivity and embodiment which are core abilities for serving as world models. This survey also covers both open-sourced models [ma2024latte, wang2025lavie, xu2024easyanimate, si2025repvideo, chen2024videocrafter2, guo2023animatediff, hacohen2024ltxvideo, chen2023videocrafter1, zheng2024opensora, zhang2024show1, lin2024opensoraplan, qing2024higen, wang2023modelscope, henschel2025streamingt2v, sun2024dimensionx, kondratyuk2023videopoet, zhou2024allegro] and commercial APIs [hailuo24minmax, klingai2024kling, gen223, pika23, girdhar2023emuvideo]. For instance, CogVideo [hong2022cogvideo] is widely regarded as an early open-source, large-scale pretrained T2V model that played a pivotal role in initiating the development of video foundation models. The release of Kling 1.0 [klingai2024kling], built on a DiT backbone, further advanced generative AI toward longer duration and greater visual quality.
@@ 4.2.1 Foundation Models with Dynamic-Aware World Knowledge
Recent research builds on this momentum, consolidating video generation as the tool for controllable world models. These models [wan2025wan, peng2025opensora2, genmo2024mochi1, xu2024easyanimate, gao2025seedance, kong2024hunyuanvideo, sora, yang2024cogvideox, klingai2024kling, imaginegrok, wan22, openai2025sora2, luma2025ray3, team2025kling, bai2025semanticgen] commonly adopt DiT-based [peebles2023dit] or some of them leverage [lin2025apt2, yin2025causvid, huang2025selfforcing, ji2025memflow, nova25amazonaws, chen2025skyreels, teng2025magi] hybrid architectures that combine diffusion models with causal, frame-by-frame autoregressive mechanisms, achieving a fast, few-step distilled diffusion process to generate each frame. For instance, models such as MAGI‑1 [teng2025magi] and CausVid [yin2025causvid] extend Diffusion Forcing [chen2024diffusionforcing] for causal video generation. Self Forcing [huang2025selfforcing] follows this paradigm but removes exposure bias by training with full autoregressive rollout that matches the true inference distribution, enabling efficient video‑level supervision and semi-real‑time generation. And VMoBA [wu2025vmoba] introduces a mixture-of-block attention mechanism to enhance spatiotemporal representation learning in video diffusion models, achieving improved generation quality and efficiency.
@@ 4.2.2 Video Generation as World Model in General Scene
In this section, we introduce methods that transition from video generation to world modeling in general-purpose scenarios, with representative methods summarized in Table 3. The term general follows its definition in the broader video generation, referring to settings that are not tailored to any specific downstream application. Our discussion focuses on two main groups of methods: (1) Approaches with spatial condition [sketchvideo, das, bian2025gsdit, li2025pisa, xue2025phyt2v, wang2025wisa, wang2025multishotmaster, li2025vfxmaster, huang2025unityvideo], which incorporate various priors such as Geometry Conditions, 3D Priors, and Physical Priors. (2) Approaches with navigation mode [fu20243dtrajmaster, he2025cameractrl2, ren2025gen3c, hou2024camtrol, aether, hong2024slowfast, xiang2024pandora, tong20254vwm], which condition generation on forms of controllable input, including action, trajectory, camera motion, and text instruction.
@@
Physical Prior World Model. Physical prior world models [li2025pisa, wang2025wisa, xue2025phyt2v, zhao2025synthetic, ji2025physmaster] aim to enhance the physical plausibility of video generation by embedding explicit physical laws and principles into the modeling process. For instance, WISA [wang2025wisa] incorporates physical formulas and principles directly as embeddings. PhyT2V [xue2025phyt2v] leverages the physical reasoning capabilities of large language models (LLMs) to guide and improve the physical consistency of video generation models.
@@
Instruction Navigation World Model. Instruction Navigation World Models [aether, xiang2024pandora, hong2024slowfast, li2024zerohsi, wei2025univideo] utilize text-based instructions, but unlike conventional text prompts that describe background elements, such as scene type, weather, or task themes, these instructions instead guide the dynamic aspects of video generation. Specifically, they control motion-related factors of either a single or multiple subjects in the video, or the perspective dynamics of a first-person agent, which can be a human, robot, or camera. For instance, models like Pandora [xiang2024pandora] and SlowFast-Gen [hong2024slowfast] can execute directional commands such as “turn left” or “turn right.” In addition, SlowFast-Gen [hong2024slowfast] supports perspective-based actions such as zooming in and out from a navigational viewpoint.